\begin{thebibliography}{10}
\expandafter\ifx\csname url\endcsname\relax
  \def\url#1{\texttt{#1}}\fi
\expandafter\ifx\csname urlprefix\endcsname\relax\def\urlprefix{URL }\fi
\expandafter\ifx\csname href\endcsname\relax
  \def\href#1#2{#2} \def\path#1{#1}\fi

\bibitem{lin1976oxygen}
S.~Lin, Oxygen diffusion in a spherical cell with nonlinear oxygen uptake
  kinetics, Journal of Theoretical Biology 60~(2) (1976) 449--457.

\bibitem{chandrasekhar1939an}
S.~Chandrasekhar, 1939an introduction to the study of stellar structure.

\bibitem{keller1955electrohydrodynamics}
J.~Keller, Electrohydrodynamics. i. the equilibrium of a charged gas in a
  container, Tech. rep., New York Univ., New York. Inst. of Mathematical
  Sciences (1955).

\bibitem{fermi1927metodo}
E.~Fermi, Un metodo statistico per la determinazione di alcune priorieta
  dell'atome, Rend. Accad. Naz. Lincei 6~(602-607) (1927) 32.

\bibitem{Bobisud1990}
L.~Bobisud, Existence of solutions for nonlinear singular boundary value
  problems, Applicable Analysis 35~(1-4) (1990) 43--57.

\bibitem{dunninger1986existence}
D.~Dunninger, J.~Kurtz, Existence of solutions for some nonlinear singular
  boundary value problem, Journal of Mathematical Analysis and Applications
  115~(2) (1986) 396--405.

\bibitem{pandey2011constructive}
R.~Pandey, A.~Verma, On a constructive approach for derivative-dependent
  singular boundary value problems, International Journal 2011.

\bibitem{ascher1995numerical}
U.~Ascher, R.~Mattheij, R.~Russell, Numerical solution of boundary value
  problems for ordinary differential equations (1995).

\bibitem{mohanty2005family}
R.~Mohanty, A family of variable mesh methods for the estimates of (du/dr) and
  solution of non-linear two point boundary value problems with singularity,
  Journal of computational and applied mathematics 182~(1) (2005) 173--187.

\bibitem{ravi2004higher}
A.~Ravi~Kanth, Y.~Reddy, Higher order finite difference method for a class of
  singular boundary value problems, Applied mathematics and computation 155~(1)
  (2004) 249--258.

\bibitem{ravi2005cubic}
A.~Ravi~Kanth, Y.~Reddy, Cubic spline for a class of singular two-point
  boundary value problems, Applied mathematics and computation 170~(2) (2005)
  733--740.

\bibitem{m2003decomposition}
M.~Inc, D.~Evans, The decomposition method for solving of a class of singular
  two-point boundary value problems, International journal of computer
  mathematics 80~(7) (2003) 869--882.

\bibitem{khuri2010novel}
S.~Khuri, A.~Sayfy, A novel approach for the solution of a class of singular
  boundary value problems arising in physiology, Mathematical and Computer
  Modelling 52~(3) (2010) 626--636.

\bibitem{ebaid2011new}
A.~Ebaid, A new analytical and numerical treatment for singular two-point
  boundary value problems via the adomian decomposition method, Journal of
  computational and applied mathematics 235~(8) (2011) 1914--1924.

\bibitem{ramos2008series}
J.~Ramos, Series approach to the Lane--Emden equation and comparison with the
  homotopy perturbation method, Chaos, Solitons \& Fractals 38~(2) (2008)
  400--408.

\bibitem{bataineh2009homotopy}
A.~Bataineh, M.~Noorani, I.~Hashim, Homotopy analysis method for singular ivps
  of Emden--Fowler type, Communications in Nonlinear Science and Numerical
  Simulation 14~(4) (2009) 1121--1131.

\bibitem{das2016algorithm}
N.~Das, R.~Singh, A.M. Wazwaz, J.~Kumar, An algorithm based on the variational
  iteration technique for the bratu-type and the Lane--Emden problems, Journal
  of Mathematical Chemistry 54~(2) (2016) 527--551.

\bibitem{singh2017optimal}
R.~Singh, N.~Das, J.~Kumar, The optimal modified variational iteration method
  for the Lane--Emden equations with neumann and robin boundary conditions, The
  European Physical Journal Plus 132~(6) (2017) 251.

\bibitem{Adomian1994a}
G.~Adomian, Solving Frontier Problems of Physics: The Decomposition Methoc [ie
  Method], Kluwer Academic Publishers, 1994.

\bibitem{adomian1983inversion}
G.~Adomian, R.~Rach, Inversion of nonlinear stochastic operators, Journal of
  Mathematical Analysis and Applications 91~(1) (1983) 39--46.

\bibitem{adomian1993}
G.~Adomian, R.~Rach, A new algorithm for matching boundary conditions in
  decomposition solutions, Applied mathematics and computation 57~(1) (1993)
  61--68.

\bibitem{adomian1994}
G.~Adomian, R.~Rach, Modified decomposition solution of linear and nonlinear
  boundary-value problems, Nonlinear Analysis: Theory, Methods \& Applications
  23~(5) (1994) 615--619.

\bibitem{wazwaz2000approximate}
A.~Wazwaz, Approximate solutions to boundary value problems of higher order by
  the modified decomposition method, Computers \& Mathematics with Applications
  40~(6-7) (2000) 679--691.

\bibitem{wazwaz2001reliable}
A.~Wazwaz, A reliable algorithm for obtaining positive solutions for nonlinear
  boundary value problems, Computers \& Mathematics with Applications
  41~(10-11) (2001) 1237--1244.

\bibitem{benabidallah2004application}
M.~Benabidallah, Y.~Cherruault, Application of the adomian method for solving a
  class of boundary problems, Kybernetes 33~(1) (2004) 118--132.

\bibitem{jang2008two}
B.~Jang, Two-point boundary value problems by the extended adomian
  decomposition method, Journal of Computational and Applied Mathematics
  219~(1) (2008) 253--262.

\bibitem{singh2013solving}
R.~Singh, J.~Kumar, Solving a class of singular two-point boundary value
  problems using new modified decomposition method, ISRN Computational
  Mathematics 2013~(Article ID 262863,) (2013) 11--pages.

\bibitem{singh2014efficient}
R.~Singh, J.~Kumar, An efficient numerical technique for the solution of
  nonlinear singular boundary value problems, Computer Physics Communications
  185~(4) (2014) 1282--1289.

\bibitem{singh2013numerical}
R.~Singh, J.~Kumar, G.~Nelakanti, Numerical solution of singular boundary value
  problems using Green's function and improved decomposition method, Journal of
  Applied Mathematics and Computing 43~(1-2) (2013) 409--425.

\bibitem{singh2014approximate}
R.~Singh, J.~Kumar, G.~Nelakanti, Approximate series solution of singular
  boundary value problems with derivative dependence using Green's function
  technique, Computational and Applied Mathematics 33~(2) (2014) 451--467.

\bibitem{singh2014adomian}
R.~Singh, J.~Kumar, The Adomian decomposition method with Green's function for
  solving nonlinear singular boundary value problems, Journal of Applied
  Mathematics and Computing 44~(1-2) (2014) 397--416.

\bibitem{singh2016efficient}
R.~Singh, A.-M. Wazwaz, J.~Kumar, An efficient semi-numerical technique for
  solving nonlinear singular boundary value problems arising in various
  physical models, International Journal of Computer Mathematics 93~(8) (2016)
  1330--1346.

\end{thebibliography}

\documentclass[12pt]{article}
\usepackage{graphicx,epsfig}
\usepackage{color}
\usepackage{amsmath,amssymb,mathrsfs,amsthm}
\usepackage{amsfonts}
\usepackage{multicol}
 \usepackage{graphicx}
\usepackage{color}
\usepackage{pdfcolmk}
\usepackage[nice]{nicefrac}
\usepackage{amsthm}
\usepackage{latexsym}
\usepackage{dsfont}
\usepackage{setspace}
\usepackage[a4paper, tmargin=1.5in, bmargin=1in, lmargin=1in, rmargin=1in, headheight=13.6pt]{geometry}
\usepackage[sort&compress,numbers]{natbib}
\usepackage[font=small,labelfont=bf,labelsep=space]{caption}
\makeatletter

\date{}

\def\@citex[#1]#2{\if@filesw\immediate\write\@auxout{\string\citation{#2}}\fi
  \def\@citea{}\@cite{\@for\@citeb:=#2\do
    {\@citea\def\@citea{,\linebreak[0]\hskip0pt plus .2em}%
      \@ifundefined{b@\@citeb}%
    {{\bf ?}\@warning{Citation `\@citeb' on page \thepage\space undefined}}%
      \hbox{\csname b@\@citeb\endcsname}}}{#1}}

\newtheorem{example}{Example}[section]

\newtheorem{rule-def}[theorem]{Rule}
\numberwithin{equation}{section}
\makeatletter

\begin{document}
\title{Adomian decomposition method for solving derivative-dependent doubly singular boundary value problems}
\author{Randhir Singh \thanks{Corresponding author. E-mail:randhir.math@gmail.com}\\
$\rm $\small {Department of Mathematics}\\
\small {Birla Institute of Technology Mesra,
 Ranchi-835215, India}} \maketitle{}
\begin{abstract}
 \noindent In this work, we apply Adomian decomposition method for solving nonlinear derivative-dependent doubly singular boundary value problems: $(py')'= qf(x,y,y')$. This method is based on the  modification of  ADM and new two-fold integral operator. The approximate  solution is obtained in the form of series with easily determinable components. The effectiveness  of the proposed approach is examined  by considering three examples and numerical results are compared with known results.
\end{abstract}
Keyword: Doubly Singular Problems;  Adomian Decomposition Method; Adomian Polynomials; Two-Point Boundary Value Problems; Derivative Dependent.
\section{Introduction}
Singular boundary value problems for ordinary differential equations arise very
frequently in many branches of applied mathematics and physics such as gas dynamics, chemical reactions, nuclear physics,  atomic structures, atomic calculations, and study of positive radial solutions of nonlinear elliptic equations and physiological studies  \cite{lin1976oxygen,chandrasekhar1939an,keller1955electrohydrodynamics,fermi1927metodo}.  Consider the following class of nonlinear  singular boundary value problems
 \begin{eqnarray}\label{sec1:1}
 (p(x)y'(x))'= q(x)f(x,y(x),y'(x)),~~0<x\leq 1,
\end{eqnarray}
with boundary conditions
\begin{align}\label{sec1:0}
y(0)=\eta_1,~\alpha_1y(1)+\beta_1y'(1)=\gamma_1,
\end{align}
where $\alpha_1>0$, $\beta_1\geq0$, $\gamma_1$ and  $\eta_1$ are any finite constants. The condition $p(0)=0$ characterizes that the problem (\ref{sec1:1}) is singular and if $q(x)$ is
allowed to be discontinuous at $x=0$ then the problem (\ref{sec1:1}) is  called doubly singular
\cite{Bobisud1990}. 

The study of singular boundary value problems  has attracted the attention of several researchers \cite{Bobisud1990,dunninger1986existence}. In \cite{dunninger1986existence}, for continuous $q(x)$ and in \cite{Bobisud1990}, for $q(x)$ not necessarily continuous at $x=0$ the existence as well as uniqueness of solution of equation \eqref{sec1:1} was discussed. The main difficulty of problem \eqref{sec1:1} is that  the singularity behavior  occurs at $x=0$. The establish of existence and uniqueness of solution of the singular problems \eqref{sec1:1}-\eqref{sec1:0} is presented in \cite{pandey2011constructive}.

Finding the solution of nonlinear SBVP \eqref{sec1:1}, most of  the methods designed for nonsingular boundary value problems suffer from a loss of accuracy or may even fail to converge \cite{ascher1995numerical}, because of the singular behavior  at $x=0$.  However, the finite difference method can be used to solve linear singular two point BVPs, but it may be difficult to solve nonlinear singular two-point boundary value problems. Moreover, the finite difference method requires some major modifications that include the use of some root-finding technique while solving nonlinear singular two-point BVPs.

In \cite{mohanty2005family}, the author has discussed existence and uniqueness of solutions of singular equation  $y''=f(x,y,y')$ and presented variable mesh methods for numerically solving such problems.  The higher order finite difference and cubic spline methods are carried out in \cite{ravi2004higher,ravi2005cubic} for the singular BVP $y''+(\nicefrac{k}{x}) y'=r(x)-q(x)y$, where $k$ is any constant.  Recently, some newly developed approximate methods have also been employed to solve special case of  SBVPs \eqref{sec1:1}. For example, Adomian decomposition method \cite{m2003decomposition,khuri2010novel,ebaid2011new}, the Homotopy perturbation method
 \cite{ramos2008series}, homotopy analysis method \cite{bataineh2009homotopy}  and variational iteration method \cite{das2016algorithm,singh2017optimal}.

\subsection{Standard Adomian decomposition method }
Adomian decomposition method has  received a lot of attention because it allows solution of both linear and nonlinear  ordinary differential, partial differential and integral equations. Adomian \cite{Adomian1994a} asserted that the ADM provides an efficient and computationally worthy method for generating approximate series solution for the large class of functional equations.
According to standard ADM    equation \eqref{sec1:1} can be rewritten as
\begin{align}\label{sec1:2}
Ly=Ny,
\end{align}
where $L\equiv\frac{d^2}{dx^2}$ is  linear  derivative operator;
 $Ny=-\nicefrac{p'}{p}y'+\nicefrac{q}{p}f(x,y,y')$ represents the nonlinear term.
The inverse  operator of  $L$ is defined  as
\begin{eqnarray}
L^{-1}(\cdot)=\int\limits_{0}^{x}\int\limits_{0}^{x}(\cdot) dx dx.
\end{eqnarray}
Operating the inverse linear operator  $L^{-1}(\cdot)$ on  both the sides
of \eqref{sec1:2}, we obtain
\begin{align}\label{sec1:3}
 y=y(0)+xy'(0)+L^{-1}Ny
\end{align}
Next, we  decompose the solution $y$ and the nonlinear function $Ny$  by an
infinite series as
\begin{align}\label{sec1:4}
y=\sum_{n=0}^{\infty}y_n,~~Ny=\sum_{n=0}^{\infty}A_n
\end{align}
where $A_n$ are Adomian polynomials that can be constructed for various classes of nonlinear functions  with the  formula given by Adomian and Rach \cite{adomian1983inversion}
\begin{eqnarray}\label{sec1:5}
A_n=\displaystyle\frac{1}{n!}\frac{d^n}{d
\lambda^n}\left[N\left(\sum_{k=0}^{\infty} y_k\lambda^k
\right)\right]_{\lambda=0},~~~n=0,1,2,...
\end{eqnarray}
Substituting  the series \eqref{sec1:4} into \eqref{sec1:3}, we obtain
\begin{align}\label{sec1:6}
\sum_{n=0}^{\infty}y_n=y(0)+xy'(0)+L^{-1}\sum_{n=0}^{\infty}A_n.
\end{align}
From the equation \eqref{sec1:6}, the ADM admits the following recursive  scheme
\begin{eqnarray}\label{e0}
\left.
  \begin{array}{ll}
  y_0=y(0)+x y'(0),\\
 y_{k+1}= L^{-1}A_{k},~~~~k\geq0.
\end{array}
\right\}
\end{eqnarray}
that will lead to the complete determination of components $y_n$ and hence the $n$-term truncated  is abstained as
\begin{eqnarray}
\psi_n(x)=\sum_{m=0}^{n-1} y_m(x),
\end{eqnarray}
The ADM has been applied to solve nonlinear boundary value problems for ordinary differential equations by many researchers   \cite{adomian1993,adomian1994,wazwaz2000approximate,
  wazwaz2001reliable,m2003decomposition,benabidallah2004application,jang2008two,khuri2010novel,ebaid2011new,singh2013solving}.
Solving nonlinear boundary value problems using standard ADM or modified ADM is always a computationally involved task as it requires the computation of undetermined coefficients in a sequence of nonlinear algebraic equations  which increases the computational work, (see \cite{benabidallah2004application,m2003decomposition,wazwaz2000approximate,wazwaz2001reliable}).

In order to avoid solving such nonlinear algebraic equations for nonlinear two-point  boundary value problems with derivative dependent source function, we use ADM which does not require any addition computational work for unknown constant based on the work \cite{singh2013solving,singh2014efficient,singh2013numerical,singh2014approximate,singh2014adomian,singh2016efficient}.

\section{Modified Adomian decomposition  method}
In this section, we apply improved ADM based on two-fold integral operator  for solving   nonlinear singular boundary value  problems. We  rewrite equation  \eqref{sec1:1} as:
\begin{align}\label{sec2:1}
Ly(x)=q(x)f(x,y,y'),
\end{align}
where $Ly(x)\equiv (p(x)y'(x))'$ is the linear differential operator to be inverted. Two-fold integral operator  $ L^{-1}(\cdot)$ regarded as the inverse operator of $L(.)$ is  proposed as
\begin{eqnarray}
L^{-1}(\cdot)\equiv\int\limits_{0}^{x}\frac{1}{p(s)}\int\limits_{s}^{1}(\cdot)dxds.
\end{eqnarray}
We operate $L^{-1}(\cdot)$ on  the left hand side of  \eqref{sec2:1} and use  $y(0)=\eta_1$ yields
\begin{align}\label{sec2:2}
\nonumber L^{-1}[(p(x)y'(x))']&=\int\limits_{0}^{x}\frac{1}{p(s)}\int\limits_{s}^{1}(p(x)y'(x))'dx
ds=\int\limits_{0}^{x}\frac{1}{p(s)}(p(1)y'(1)-p(s)y'(s))ds,\\
&=c_1\int\limits_{0}^{x}\frac{1}{p(s)}ds-y(x)+\eta_1=c_1\int\limits_{0}^{x}\frac{1}{p(s)}ds-y(x)+\eta_1.
\end{align}
We  again operate  $L^{-1}(.)$ on both sides of  \eqref{sec2:1} and use \eqref{sec2:2},  gives
\begin{eqnarray}\label{sec2:3}
y(x)=\eta+c_1\int\limits_{0}^{x}\frac{1}{p(s)}ds-\int\limits_{0}^{x}\frac{1}{p(s)}\int\limits_{s}^{1}q(x)f(x,y,y')dxds.
\end{eqnarray}
For simplicity, we set $$h(x)= \int\limits_{0}^{x}\frac{1}{p(s)}ds,~~
[L^{-1}(.)]_{x=1}=\int\limits_{0}^{1}\frac{1}{p(s)}\int\limits_{s}^{1}(.)dxds,~~~~~h'(1)=\frac{1}{p(1)}.$$
Then the equation \eqref{sec2:3} may be written as
\begin{eqnarray}\label{sec2:4}
y(x)=\eta_1+c_{1}h -[L^{-1}q f(x,y,y')].
\end{eqnarray}
 To eliminate  $c_1$ from \eqref{sec2:4},  we impose $\alpha_1y(1)+\beta_1y'(1)=\gamma_1$, and we obtain
 \begin{eqnarray}\label{sec2:5}
c_1=\frac{(\gamma_1-\alpha_1\eta)}{\alpha_1h(1)+\beta_1 h'(1)}+\frac{\alpha_1}{\alpha_1h(1)+\beta_1 h'(1)}
\big[L^{-1}q f(x,y,y')\big]_{x=1}.
\end{eqnarray}
Substituting the value of $c_1$ into \eqref{sec2:4}, we obtain
\begin{align}\label{sec2:6}
\displaystyle y(x)=\eta_1+\frac{(\gamma_1-\alpha_1\eta_1) h}{\alpha_1h(1)+\beta_1 h'(1)}+\frac{\alpha_1 h}{\alpha_1h(1)+\beta_1 h'(1)}
\bigg[L^{-1}q f(x,y,y')\bigg]_{x=1}-\big[L^{-1}q f(x,y,y')\big].
\end{align}
Substituting the series defined in \eqref{sec1:4} into  \eqref{sec2:6} gives
\begin{equation}\label{sec2:eq9}
\sum_{n=0}^{\infty}y_n=\eta_1+\frac{(\gamma_1-\alpha_1\eta_1)h}{\alpha_1h(1)+\beta_1 h'(1)}+\frac{\alpha_1 h}{\alpha_1h(1)+\beta_1 h'(1)}
\big[L^{-1}q\sum_{n=0}^{\infty} A_n\big]_{x=1}-\bigg[L^{-1}q\sum_{n=0}^{\infty} A_n \bigg].
\end{equation}
Comparing  the both sides of  \eqref{sec2:eq9}, we have
\begin{eqnarray}\label{f2}
\left.
  \begin{array}{ll}
\displaystyle    y_0=\eta_1,\\
\displaystyle y_1=\frac{(\gamma_1-\alpha_1\eta)h}{\alpha_1h(1)+\beta_1 h'(1)}+\frac{\alpha_1h}{\alpha_1h(1)+\beta_1 h'(1)}\big[L^{-1}qA_{0}\big]_{x=1}-\big[L^{-1}qA_{0} \big],\\
\vdots\\
\displaystyle y_{n+1}=\frac{\alpha_1h(x)}{\alpha_1h(1)+b
h'(1)}\big[L^{-1}q A_{n}\big]_{x=1}-\big[L^{-1}q A_{n}\big],~~~~n\geq1.
  \end{array}
\right\}
\end{eqnarray}
The  recursive scheme  \eqref{f2}
gives the complete determination of solution components $y_n$ of solution $y$ and hence the approximate  series solution $\psi_n$ can be obtained as
\begin{eqnarray}\label{sec2:eq10}
\psi_n=\sum_{m=0}^{n-1} y_m.
\end{eqnarray}

\section{Numerical illustrations and discussions}
In order to demonstrate the effectiveness and efficaciously of the proposed method, we have considered three complicated nonlinear singular examples. All the numerical results obtained by proposed ADM are compared with known results. Now,  we denote error functions as $E_n(x)=|\psi_n(x)-y(x)|$  and the maximum absolute errors as
 \begin{align}
E^{n}=\max_{0<x\leq1}E_n(x).
\end{align}

\begin{example}\label{sec3:ex1}
{\rm Consider the nonlinear  boundary value problem
\begin{eqnarray}\label{prob_1}
\left.
  \begin{array}{ll}
(x^{\alpha}y')'= \beta x^{\alpha+\beta-2}e^{y}\big( -x y'-\alpha-\beta+1 ),~~~~0<x\leq 1,\\
y(0)=\ln\frac{1}{4}~~~y(1)=\ln\frac{1}{5},
\end{array}
\right\}
\end{eqnarray}
with exact solution $y(x)=\ln(\frac{1}{4+x^{\beta}})$.}
\end{example}
Applying \eqref{f2} to the equation \eqref{prob_1}, with  $\alpha_1=1$, $\beta_1=0$,
 $\gamma_1=-\ln(5)$ and $\eta=-\ln(4)$, we have
\begin{align}\label{sec3:s1}
\left.
  \begin{array}{ll}
 y_0=-\ln(4),\\
 y_1=(\ln4-\ln5)x^{1-\alpha}+x^{1-\alpha}\int\limits_{0}^{1}x^{-\alpha}\int\limits_{x}^{1}x^{\alpha +\beta -2}A_{0}dxdx -\int\limits_{0}^{x}x^{-\alpha}\int\limits_{x}^{1}x^{\alpha +\beta -2}A_{0}dxdx,\\
y_{n+1}=x^{1-\alpha}\int\limits_{0}^{1}x^{-\alpha}\int\limits_{x}^{1}x^{\alpha +\beta -2}A_{n}dxdx-
\int\limits_{0}^{x}x^{-\alpha}\int\limits_{x}^{1}x^{\alpha +\beta -2}A_{n}dxdx,~~n\geq1.
  \end{array}
\right\}
\end{align}
The Adomian polynomials  for  $f(x,y,y')=-\beta \big(xe^{y}y'+e^{y}(\alpha+\beta-1) )$ with $y_0=-\ln(4)$ are obtained as:
\begin{align}\label{sec3:a1}
\left.
 \begin{array}{ll}
 A_0=-\beta e^{y_0}\big(x y'_0+(\alpha+\beta-1)\big)  \\
 A_1=-\beta e^{y_0}\big(x y'_1+y_1(\alpha +\beta -1) \big),\\
 A_2=-\beta e^{y_0}\big(x y'_2+y_1 y'_1)+( y_2+\nicefrac{y_1^2}{2} )(\alpha -\beta +1)\big),\\
 A_3=-\beta  e^y\big(x \left( y'_3+y'_2y_1+y'_1\left(\nicefrac{y_1^2}{2}+y_2\right)\right)+\left( y_3+y_1 y_2 +\nicefrac{y_1^3}{6} \right)(\alpha +\beta -1) \big),\\
\ldots
\end{array}
\right\}
\end{align}

For~$\alpha=0.5$,~$\beta=1$: Using \eqref{sec3:s1} and \eqref{sec3:a1},  we obtain
\begin{align*}
y_0&=-1.38629436,\\
y_1&=0.0268564x^{0.5}-0.25 x,\\
y_2&=-0.0267739x^{0.5}-0.00447607 x^{1.5}+0.03125 x^{2},\\
y_3&=-0.0003279x^{0.5}+0.0044623x^{1.5}-0.000045079x^{2}+0.00111902x^{2.5}-0.0052083x^{3},\\
y_4&=0.00025327x^{0.5}+0.0000546545 x^{1.5}+0.0000898816 x^{2}-0.0011159 x^{2.5}\\&+0.0000212874 x^{3}-0.0002797x^{3.5}+0.000976563 x^{4},\\
\ldots
\end{align*}
For~$\alpha=0.5$,~$\beta=3.5$: Using \eqref{sec3:s1} and \eqref{sec3:a1},  we obtain
\begin{align*}
y_0&=-1.38629436,\\
y_1&=0.+0.0268564 x^{0.5}-0.25 x^{3.5},\\
y_2&=0.-0.0253752 x^{0.5}-0.00587485 x^{4}+0.03125 x^{7},\\
y_3&=-0.00174107 x^{0.5}+0.00555081x^{4}-0.000070123x^{4.5}+0.00146871x^{7.5}-0.00520833x^{10.5},\\
y_4&=0.000230726 x^{0.5}+0.000380859 x^{4}+0.000132511 x^{4.5}-5.6497931825\times 10^{-7} x^{5}\\&-0.0013877 x^{7.5}+0.0000347878 x^{8}-0.000367178 x^{11}+0.000976563 x^{14},\\
\ldots
\end{align*}
The maximum absolute error $E^{(n)}$ for $n=5,8$, and $10$ are listed in Table \ref{tab1} and \ref{tab2} for different values of $\alpha$ and $\beta$.
\begin{table}[htbp]\caption{Maximum error of Example~\ref{sec3:ex1}, when $\beta=1$.}
\centering
\vspace{-.3cm}
\renewcommand{\arraystretch}{1.0}
\setlength{\tabcolsep}{0.25in}
\begin{tabular}{c|c cc}
\cline{1-4}
$\alpha$ &  $E^{(5)}$               &$E^{(8)}$               & $E^{(10)}$ \\
\cline{1-4}
0.25    &1.11205$\times 10^{-5}$  &2.30301$\times 10^{-8}$   &3.54171$\times 10^{-10}$ \\
0.5     &1.52386$\times 10^{-5}$  &2.68725$\times 10^{-8}$   &6.11240$\times 10^{-10}$ \\
0.75    &1.62907$\times 10^{-5}$  &4.91114$\times 10^{-8}$   &9.22449$\times 10^{-10}$ \\
\hline
\end{tabular}
\label{tab1}
\end{table}
\begin{table}[htbp]\caption{Maximum error of Example \ref{sec3:ex1}, when $\beta=3.5$.}
\centering
\vspace{-.3cm}
\renewcommand{\arraystretch}{1.0}
\setlength{\tabcolsep}{0.25in}
\begin{tabular}{c|c cc}
\cline{1-4}
$\alpha$ &  $E^{(5)}$               &$E^{(8)}$               & $E^{(10)}$ \\
\cline{1-4}
0.25    &1.50942$\times 10^{-5}$  &5.61969$\times 10^{-8}$   &1.12729$\times 10^{-9}$ \\
0.5     &1.58204$\times 10^{-5}$  &6.53078$\times 10^{-8}$   &1.32773$\times 10^{-9}$ \\
0.75    &2.91795$\times 10^{-5}$  &6.92627$\times 10^{-8}$   &1.38617$\times 10^{-9}$ \\
\hline
\end{tabular}
\label{tab2}
\end{table}

\begin{example}\label{sec3:ex2}
{\rm Consider the nonlinear singular boundary value problem
\begin{eqnarray}\label{prob_2}
\left.
  \begin{array}{ll}
(x^{\alpha}y')'=  x^{\alpha-1}e^{y}\big( -x y'-\alpha ),~~~~0<x\leq 1,\\
y(0)=\ln\frac{1}{2},~~y(1)=\ln\frac{1}{3},
\end{array}
\right\}
\end{eqnarray}
and exact solution is $y(x)=\ln(\frac{1}{2+x})$.}
\end{example}
Applying \eqref{f2} to the equation to the equation \eqref{prob_2}, with $\alpha_1=1$, $\beta_1=0$, $\gamma_1=-\ln3$ and $\eta_1=-\ln2$, we obtain
\begin{align}\label{sec3:s2}
\left.
  \begin{array}{ll}
 y_0=-\ln2,\\
 y_1=(\ln2-\ln3)x^{1-\alpha}+x^{1-\alpha}\int\limits_{0}^{1}x^{-\alpha}\int\limits_{x}^{1}x^{\alpha -1}A_{0}dxdx -\int\limits_{0}^{x}x^{-\alpha}\int\limits_{x}^{1}x^{\alpha -1}A_{0}dxdx,\\
 y_{n+1}=x^{1-\alpha}\int\limits_{0}^{1}x^{-\alpha}\int\limits_{x}^{1}x^{\alpha- 1}A_{n}dxdx-
\int\limits_{0}^{x}x^{-\alpha}\int\limits_{x}^{1}x^{\alpha -1}A_{n}dxdx,~~n\geq1.
  \end{array}
\right\}
\end{align}
The Adomian polynomials   for  $f(x,y,y')=-\big( x e^{y} y'+e^{y}\alpha )$ with $y_0=-\ln(2)$ are obtained as
\begin{align}\label{sec3:a2}
\left.
 \begin{array}{ll}
 A_0=- e^{y_0}\big(x y'_0+\alpha\big)  \\
 A_1=- e^{y_0}\big(x y'_1+y_1\alpha  \big),\\
 A_2=-e^{y_0}\big(x y'_2+y_1 y'_1)+( y_2+\nicefrac{y_1^2}{2} )\alpha \big),\\
 A_3=- e^{y_0}\big(x \left( y'_3+y'_2y_1+y'_1\left(\nicefrac{y_1^2}{2}+y_2\right)\right)+\left( y_3+y_1 y_2 +\nicefrac{y_1^3}{6} \right)\alpha  \big),\\
\ldots
\end{array}
\right\}
\end{align}
For~$\alpha=0.5$:  Using \eqref{sec3:s2} and \eqref{sec3:a2},  we obtain 
\begin{align*}
y_0&=-0.693147,\\
y_1&=0.0945349x^{0.5}-0.5x,\\
y_2&=-0.0934884 x^{0.5}-0.0315116 x^{1.5}+0.125 x^{2},\\
y_3&=-0.00413483 x^{0.5}+0.0311628 x^{1.5}-0.00111711 x^{2}+0.0157558 x^{2.5}-0.0416667 x^{3},\\
y_4&=0.00321966 x^{0.5}+0.00137828 x^{1.5}+0.00220948 x^{2}-0.0156096 x^{2.5}+0.00105504 x^{3}\\&-0.00787791 x^{3.5}+0.015625 x^{4},\\
\ldots
\end{align*}
For different values of $\alpha$, the maximum absolute error $E^{(n)}$, for $n=5,8$ and $10$, are given in  Table \ref{tab3}.  It can be noted that  when $n$  increases the maximum error decreases.
\begin{table}[htbp]\caption{Maximum error of Example~\ref{sec3:ex2}}
\centering
\vspace{-.3cm}
\renewcommand{\arraystretch}{1.0}
\setlength{\tabcolsep}{0.25in}
\begin{tabular}{c|c cc}
\cline{1-4}
$\alpha$ &  $E^{(5)}$               &$E^{(8)}$               & $E^{(10)}$ \\
\cline{1-4}
0.25    &8.25847$\times 10^{-5}$  &8.22973$\times 10^{-7}$   &3.94246$\times 10^{-8}$ \\
0.5     &8.41014$\times 10^{-5}$  &1.68118$\times 10^{-6}$   &4.29665$\times 10^{-8}$ \\
0.75    &3.17953$\times 10^{-5}$  &8.27238$\times 10^{-6}$   &6.58700$\times 10^{-8}$ \\
\hline
\end{tabular}
\label{tab3}
\end{table}

\begin{example}\label{sec3:ex3}
{\rm Consider the  singular boundary value  problem
\begin{eqnarray}\label{prob_3}
\left.
  \begin{array}{ll}
    (x^{\alpha}y')'=\beta x^{\alpha+\beta-2}\left( xy'+y(\alpha+\beta-1)\right),~~0<x\leq1, \\
  y(0)=1~~~ {\rm and}~~~ y(1)=e,
\end{array}
\right\}
\end{eqnarray}
with exact solution $y(x)=e^{x^\beta}$. }
\end{example}
Applying \eqref{f2} to the equation to the equation \eqref{prob_3}, with  $\alpha_1=1$, $\beta_1=0$, $\gamma_1=e$ and $d=\eta_1$, we obtain
\begin{align}\label{sec3:s3}
\left.
  \begin{array}{ll}
    y_0=1,\\
y_1=(e-1)x^{1-\alpha}+x^{1-\alpha}\int\limits_{0}^{1}x^{-\alpha}\int\limits_{x}^{1}x^{\alpha +\beta -2}A_{0}dx dx -
\int\limits_{0}^{x}x^{-\alpha}\int\limits_{x}^{1}x^{\alpha +\beta -2}A_{0}dxdx,\\
 y_{n+1}=x^{1-\alpha}\int\limits_{0}^{1}x^{-\alpha}\int\limits_{x}^{1}x^{\alpha +\beta -2}A_{n}dxdx-
\int\limits_{0}^{x}x^{-\alpha}\int\limits_{x}^{1}x^{\alpha +\beta -2}A_{n}dxdx,~~n\geq1,
  \end{array}
\right\}
\end{align}
where $A_n=\beta  \left( xy'_n+y_n(\alpha+\beta-1)\right).$

For~$\alpha=0.5$,~$\beta=1$: Using \eqref{sec3:s3},  we obtain
\begin{align*}
y_0&=1,\\
y_1&=0.718282 x^{0.5}+x,\\
y_2&=-0.978855 x^{0.5}+0.478855 x^{1.5}+0.5 x^{2},\\
y_3&=0.294361 x^{0.5}-0.65257 x^{1.5}+0.191542 x^{2.5}+0.166667 x^{3},\\
y_4&=-0.0316058 x^{0.5}+0.196241 x^{1.5}-0.261028 x^{2.5}+0.0547262 x^{3.5}+0.0416667 x^{4},\\
\ldots
\end{align*}
 
For $\alpha=0.5$,~$\beta=2.5$:  Using \eqref{sec3:s3},  we obtain
\begin{align*}
y_0&=1,\\
y_1&=0.718282 x^{0.5}+ x^{2.5},\\
y_2&=-1.09857 x^{0.5}+0.598568 x^{3}+0.5 x^{5},\\
y_3&=0.47673 x^{0.5}-0.915473 x^{3}+0.272076 x^{5.5}+0.166667 x^{7.5},\\
y_4&=-0.107842 x^{0.5}+0.397275 x^{3}-0.416124 x^{5.5}+0.0850239 x^{8}+0.0416667 x^{10},\\
\ldots
\end{align*}
The maximum absolute error for different values of $\alpha$ and $\beta$ in Table\ref{tab4} and  Table \ref{tab5}.
 \begin{table}[htbp]\caption{Maximum error of Example~\ref{sec3:ex3},when $\beta=1$}
\centering
\vspace{-.3cm}
\renewcommand{\arraystretch}{1.0}
\setlength{\tabcolsep}{0.25in}
\begin{tabular}{c|c cc}
\cline{1-4}
$\alpha$ &  $E^{(5)}$               &$E^{(8)}$               & $E^{(10)}$ \\
\cline{1-4}
0.25    &9.47534$\times 10^{-5}$  &9.45679$\times 10^{-7}$   &2.28700$\times 10^{-9}$ \\
0.5     &6.54946$\times 10^{-5}$  &6.83854$\times 10^{-7}$   &1.04295$\times 10^{-8}$ \\
0.75    &8.17721$\times 10^{-5}$  &1.08101$\times 10^{-6}$   &1.18947$\times 10^{-8}$ \\
\hline
\end{tabular}
\label{tab4}
\end{table}
 \begin{table}[htbp]\caption{Maximum error of Example~\ref{sec3:ex3},when $\beta=2.5$}
\centering
\vspace{-.3cm}
\renewcommand{\arraystretch}{1.0}
\setlength{\tabcolsep}{0.25in}
\begin{tabular}{c|c cc}
\cline{1-4}
$\alpha$ &  $E^{(5)}$               &$E^{(8)}$               & $E^{(10)}$ \\
\cline{1-4}
0.25    &8.48845$\times 10^{-4}$  &5.11258$\times 10^{-6}$   &3.01254$\times 10^{-8}$ \\
0.5     &9.99777$\times 10^{-4}$  &5.94699$\times 10^{-6}$   &3.44033$\times 10^{-8}$ \\
0.75    &2.97716$\times 10^{-4}$  &1.53819$\times 10^{-6}$   &1.19534$\times 10^{-8}$ \\
\hline
\end{tabular}
\label{tab5}
\end{table}

\section{Conclusion}
 The Adomian decomposition method has been used  for  solving a class of nonlinear singular boundary value problems  with derivative dependent  source function, i.e., $f(x,y,y')$.  The main advantage of the approach is that it provides a direct scheme for solving the doubly singular boundary value  problems.  The method provides a reliable technique which requires less work compared to standard Adomian decomposition method. The numerical results of the examples are presented  and only a few terms are required to obtain accurate solutions.  By comparing the results with other existing methods, it has been proved that  ADM is a more powerful method for solving doubly  singular boundary value problems.


\end{document}